The peak concentration of organic aldehydes on a smoggy day reaches 0.5 ppm. If the aldehydes condense into organic aerosol with average particle diameters of 1 um and average densities of 1 g/cm3, estimate the aerosol particle concentrations resulting from the aldehydes
Answer #1

1 ppm = 1 mg/L

So, peak concentration of organic aldehydes = 0.5 mg/L

                                                                     = (0.5*10-3 g)/(1000 cm3)

                                                                     = 5*10-7 g/cm3

Given, average particle diameter = 1 μm

Average particle radius = (1/2) μm

= 0.5 μm

= 0.5*10-4 cm   (1 μm = 10-4 cm)

Average particle volume = (4/3)*π* (0.5*10-4) cm3

                                                                = 5.235*10-13 cm3

Average particle mass = Average particle volume * density

                                      = 5.235*10-13 cm3 * 1g/cm3

                                      = 5.235*10-13 g

Let there be ‘n’ aerosol particle per cm3 i.e. the aerosol particle concentration is ‘n’ particles/cm3.

So,   n*5.235*10-13 = 5*10-7

Or, n = 9.55*105 particles/cm3 (ANSWER)